OpenStudy (anonymous):

Find the slope of the tangent line to the curve 4 x^2 -2 xy + 2 y^3 = 94 at the point ( 4,3 )

OpenStudy (anonymous):

\[find \frac{ dy }{ dx } and then plugx=4,y=3\]

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

wait, i keep getting it wrong...:( @surjihayer

OpenStudy (anonymous):

is this the derivative? 8x-2y'+6yy'=0? @Hero

OpenStudy (anonymous):

\[8x-2\left( xy \prime+1y \right)+6y ^{2}y \prime =0\] \[\left( -2x+6y ^{2} \right)y \prime =-8x+2y\] \[y \prime=\frac{ 2y-8x }{ 6y ^{2}-2x }=\frac{ y-4x }{3y ^{2}-x }\] now plug the values.

OpenStudy (anonymous):

i cant read that, is it different from my answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{ d }{dx }\left( xy \right)=x \frac{ dy }{ dx}+1y\]

OpenStudy (anonymous):

\[\frac{ d }{ dx}\left( 2y ^{3} \right)=2*3 y ^{2}\frac{ dy }{dx }\]

OpenStudy (anonymous):

\[\frac{ d }{dx }\left( xy \right)=x \frac{ dy }{ dx}+1y\]

OpenStudy (anonymous):

yes